OpenStudy (anonymous):

Solve the system, using substitution. Write the solution as an ordered pair (a,b) 3a - b = 11 and 2a + 3b = 0 (3, 2) (3, -2) (-3, 2) (-3, -2) no solution

OpenStudy (anonymous):

All you have to do is simplify, and then substitute one equation onto the other i will isolate b in the first one. 3a - b = 11 -b = -3a + 11 b = 3a - 11 then enter it in the second equation 2a + 3b = 0 2a + 3(3a - 11) = 0 simplify 2a + 9a - 33 = 0 11a = 0 + 33 11a = 33 a = 33 / 11 a = 3 the i plot my a answer in the equations to find the answer for b 3a - b = 11 3(3) - b = 11 9 - b = 11 -b = 11 - 9 -b = 2 b = -2 2a + 3b = 0 2(3) + 3b = 0 6 + 3b = 0 3b = 0 - 6 3b = -6 b = -6 / 3 b = -2 The answer would be (3,-2) if a =x and b = y